Calculate the pH of a solution containing 0.1 mol of phosphoric acid per liter of solution at SATP knowing that the SATP values of pK1, pK2, and pK3 for H3PO4 are 2.1, 7.2, and 12.4, respectively.

Show your work.
Answer #1

K for HyPO,-101 -2.1 =7.9×10-3 I 0.1 x2-7.9*10-3 (0.1-x)=0 By solving above equati on we get x 0.0244 M Hox 0.0244M pHlogHo]